OpenStudy (anonymous):

Find how many quarts of 4% low fat milk should be mixed to yield 60 quarts of 3 % low fat milk. -The mixture should contain---quarts of the 4% lowfat milk. -The mixture should contain--- quarts of the 1% low fat milk. Any help???? THANK YOU:)

hero (hero):

x = 4% low fat milk y = 100% water Total amount of liquid in quarts: x + y = 60 Percentage of low fat milk equation: .04x + 0y = .03(60)

hero (hero):

So solve this system: x + y = 60 .04x + 0y = 1.8 When you solve the system you get x = 45 y = 15 So the amount of 4% low fat milk that should be mixed is 45 quarts

hero (hero):

I didn't read anything about 1% lowfat milk in the problem. So if you did not post the problem completely, you should do so.